A bird has a mass of 26 g and perches in the middle of a stretched telephone line. (a) Show that the tension in the line can be calculated using the equation T=mg/2sinθ. 


1
Expert's answer
2020-07-13T11:44:13-0400

The free-body diagram is shown on the figure:


As far as the bird doesn't move, its acceleration "a = 0". Thus, writingg the second Newton's law in a projection on the vertical axis, obtain:


"mg - T\sin\theta - T\sin \theta = 0\\
mg = 2T\sin\theta\\
T = \dfrac{mg}{2\sin\theta}"

QED